# +SSLIP: Automated Radon-assisted and Rotation-corrected identification of complex HCP slip system activity fields from DIC data

### v1

@@ 2.3.1 SSLIP Analysis with Preselected Slip System Combinations
With the set of preselected slip systems established, we proceed to explain the SSLIP analysis phase. This involves evaluating every possible combination of two slip systems from the preselected pool for each datapoint within a certain grain. Specifically, for each pair of slip systems, we apply the SSLIP methodology, utilizing the optimization equation (Section 2.1). Additionally, rotation correction will be applied, as explained below in Section 2.4.
The selection of the optimal combination for each datapoint is governed by the residual fraction, $Rf_{2D}^{L_{2}}$, which must fall below a predefined threshold to ensure accurate identification. This residual fraction is calculated as
$Rf_{2D}^{L_{2}}=\frac{\left\|\mathbf{H}^{exp}-\mathbf{H}^{theo}\right\|^{2D}}{E_{eff}},$ (10)
where $\mathbf{H}^{exp}$ represents the experimental displacement gradient tensor, $\mathbf{H}^{theo}$ denotes the theoretical displacement gradient tensor predicted by a combination of slip systems and rotation (see Section 2.4), and $E_{eff}$ is the effective strain, serving as a normalization factor.
The constraints imposed on potential combinations are outlined in Table 1.
@@
Each combination is thereby assessed based on its residual fraction (Eq. 10), rotation angle and minimum equivalent strain, according to the values specified in Table 1. These values have been chosen as they were found to work robustly for the analysis of hundreds of Zn grains, of which a few challenging cases are discussed in Section 3. Among these valid combinations, the one with the lowest sum of slip amplitudes, thereby indicating the path of least resistance and conforming to the principle of minimum energy, is selected as the definitive combination of slip system activities for that particular datapoint.
@@
Since SSLIP will be applied with only two systems at a time, the optimization problem will have fewer degrees of freedom, which allows us to simultaneously identify the local elastic rotations. Inclusion of local elastic rotations was not needed earlier for the less challenging cases of FCC and BCC, but is found to be necessary for HCP because of its large number slip systems with linearly dependent in-plane deformation gradient tensor (due to the high anisotropy and wide range of slip families of HCP). In practice, these rotations can be in the form of (full) grain rotation (e.g. due to grain boundary sliding), by compatibility effects near a grain boundary or other microstructural features [4], or simply by misalignment of the sample between (quasi-situ) measurement steps. Therefore, we add (local) rotation correction to the methodology, which will not only improve the slip system identification but will also give an extra result, i.e. the rotation field, which may be useful for the analysis.
The correction involves incorporating an additional rotation component into SSLIP, represented by a simplified rotation tensor. For small rotations, which are common in our observations, the in-plane rotation tensor can be described and simplifies as follows:
@@
Here, $\theta$ represents the rotation angle and, under the assumption that $\theta$ is small, $\sin(\theta)\approx\theta$ and $\cos(\theta)\approx 1$, allowing us to capture the in-plane rotation approximately with this simplified matrix, in displacement gradient tensor form. It is then incorporated into SSLIP as an additional slip system, for each combination that will be tried. This results in the rotation value $\theta$ at every datapoint, along with the slip system activities. In cases where the rotations would be so large that the simplified rotation tensor is a poor estimate, then the full version of $\mathbf{H}^{rot}$ in Eq. 11 can be used, at the expense that $\mathbf{H}^{rot}$ needs to be updated using the current estimate of the rotation angle, at each location and at each step in the iterative optimization routine.
@@ 2.5 Validation on Virtual Experiment
Now, the full +SSLIP methodology is applied to the challenging HCP virtual experiment, for which the results are shown in Figure 5. This figure is structured identically to those in the experimental case study below. Figure 5(a) shows the crystallographic orientation of the virtual experiment and Figure 5(b) shows again the effective strain field. Application of the Radon transfer results in 8 preselected slip systems, see Figure 2. Subsequently, application of Rotation-corrected SSLIP on all pairs of the slip systems ($N=28$) and point-wise selection of the pair with the lowest residual results in the slip activity fields as shown in Figure 5. The correct slip systems were Basal-3, PrismI-3, PyrICA-9 and PyrII-5, all of which show a single slip band, oriented as expected along their slip trace. The other 4 pre-selected, but non-active, slip systems show no activity over the full area, as expected. The error of the SSLIP analysis is represented by the residual norm fraction in Figure 5(d), $Rf_{2D,max}^{L_{2}}$ (Eq. 10), in which low values can be observed. No rotation was applied in the virtual experiment, which is reproduced by the +SSLIP algorithm as shown in Figure 5(c). The intersection of 3 slip bands in the center of the domain cannot be identified, since only 2 slip systems can be active per datapoint. Importantly, these intersections are also not falsely attributed to incorrect slip systems. In practice, the slip activity at such multi-slip intersections can easily be inferred from the neighbouring slip activity. Overall, the identification shown here is robust and accurate and shows promise for identification of complex HCP activities.
@@
Figure 5: Demonstration of the +SSLIP results on the challenging virtual experiment. (a) The EBSD map and rotated crystal shape illustrates the crystal orientation of this virtual experiment. (b) The effective strain field, $E_{eff}$, showing the deformation pattern, with 4 slip bands clearly visible. (c) The rotation field, $\gamma_{rot}$, resulting from rotation-correction by +SSLIP . (d) The residual strain fraction field, $Rf_{2D}^{L_{2}}$, resulting from +SSLIP, showing the fit between the ”experimental” and theoretical displacement gradient tensor. Below that, the slip system activity map of each slip system of each slip family (Basal, PrismI, PyrICA and PyrII) is shown. Maps with red crosses indicate that systems were not selected by Radon transform preselection. Each map also features the slip trace (red line) and direction (orange arrow), with the Schmid factor ($SF$) depicted as colored circles (colorbar below (a)). The SF is not used in the +SSLIP method.
2.6 Conclusive vs. nondiscriminatory identification
@@
At this point it may seem that the +SSLIP method is working perfectly, at least on this virtual test, but one issue has been left undiscussed, i.e. the possible non-uniqueness of the 2D part of a slip system’s displacement gradient tensor for specific grain orientations. In theory, even though each slip system has of course a unique 3D displacement gradient, the 2D part of the displacement gradient tensor could be exactly, or almost, the same for two (or more) slip systems, making those slip systems fundamentally nondiscriminatory in the SSLIP method. In practice, however, the minimization of the sum of slip activities, as used in the original SSLIP method [22], was found to be robust against slip systems with almost equal 2D displacement gradient tensors, therefore, this issue was not considered in detail before. For HCP, however, there are many potential slip systems from multiple different families (which typically cannot be excluded a-priori due to unknown CRSS values) and the occurrence of multiple highly similar 2D displacement gradient tensors was found to be problematic for grains with ’unfortunate’ grain orientation, a few examples of which are treated in Section 3. In such cases, a group of 2 or more slip systems can be defined with similar enough 2D displacement gradient tensors such that the measured slip activity can only be assigned to the group of slip systems. When slip activity is assigned to such a ’nondiscriminatory slip system group’ we call the identification ’inconclusive’, as counterpart to the regular ’conclusive’ identification where the slip activity is uniquely assigned to one or more slip systems.
@@
Therefore, in the +SSLIP method, after the Radon transform preselection, all possible nondiscriminatory slip system groups are identified from the available preselected slip systems, before executing the +SSLIP identification. Whether or not the theoretical 2D displacement gradient tensor of a preselected slip systems can be discriminated from the experimentally measured 2D displacement gradient tensor caused by activity of another slip systems depends on the absolute measurement accuracy of the slip trace angle and the projected slip direction angle, which depends on the combined accuracy of EBSD, SEM-DIC (including SEM scanning artefacts), and their mutual alignment. For our case, it was found that these two angles can have an error up to $\sim$5. Therefore, we define an angle threshold of 5 , i.e. if two preselected slip systems have both a slip trace angle and projected slip direction angle within 5of each other, they are considered nondiscriminatory. Through this procedure, the preselected slip systems are, when necessary, included in nondiscriminatory slip system groups in the +SSLIP identification. In the (pixel-wise) +SSLIP output, the slip activity assigned to a nondiscriminatory group is attributed equally to each slip system within the group. The way this works in practice will be shown in the section below (note that the virtual test case in Figure 5 is a regular example of a ’conclusive’ identification).
@@ 3 Experimental Case Study: +SSLIP applied on HCP Zinc Coating
In this section, we demonstrate the strength of the +SSLIP method on experimental SEM-DIC results of an HCP Zinc anti-corrosion coating, which will show activities from multiple slip families: Basal, PyrI and PyrII. The focus will be on challenging cases that include, .e.g., multiple ’nondiscriminatory slip system groups’. Additionally, we will showcase robust identification on grains that exhibit cross slip and diffuse slip, for which pure trace-based analysis is hardly possible [22].
In Section 3.1, we briefly detail the experimental methodology for acquiring microstructure-aligned SEM-DIC strain fields. Section 3.2 demonstrates +SSLIP a simple case of single slip and Section 3.3 shows how an inconclusive identification can be interpreted. More complex cases of multi-slip, cross slip and diffuse slip are treated in Sections 3.4 and 3.5.
@@ 3.1 Material Characterization and SEM-DIC testing
The material used in this case study is a hot-dip galvanized skin-passed Zn coated steel sample whose material code is DX54-Galvanized Iron (GI). The thickness of the Zn anti-corrosion coating is approximately 10 $\mu$m. Dogbone samples are cut using wire-EDM such that they can be deformed using a Kammrath&Weiss tensile stage mounted inside a Tescan Mira 3 SEM, at close working distance, in order to optimize the spatial resolution of SEM-DIC [32, 5]. The gauge width is 4 mm and sample thickness is $0.7$ mm. Before testing, the sample surface is carefully prepared in a single step of polishing using Oxide Polishing Suspension (OPS)-NonDry and ethanol (1:1) suspension on a Struers MD-Chem cloth for 10 minutes. This short polishing provides the optimal compromise between good surface quality and low material removal, since the Zn coating is very thin.
@@
Next, several $\sim$50 $\mu$m sized Regions of Interest (RoI’s) are chosen for characterization and in-situ SEM-DIC testing, as shown in Figure 6 for one RoI. EBSD is performed with an EDAX Digiview 2 camera in a Tescan Mira 3 SEM, with offline spherical indexing using EMSphInx for improved quality [33], resulting in an Inverse Pole Figure (IPF) map as visualized in Figure 6(a). Subsequently, we apply a fine and dense InSn SEM-DIC speckle pattern applied according to the parameters of pattern b in Table 1 of Hoefnagels et al. [34], resulting in approximately 90-100 nm sized speckles, see Figure 6(b). The specimen is then mounted in a Kammrath&Weiss micro-tensile stage, which is installed inside the Tescan Mira 3 SEM for interrupted in-situ testing, at deformation steps as shown in Figure 6(d). For SEM-DIC imaging, we use in-lens SE imaging at a 5 kV beam voltage and at 7 mm working distance. At each deformation step, a horizontal and vertical scan are captured that are combined into a single image that is corrected for SEM scanning artefacts, using a software program called ScanCorr [4]. Subsequently, DIC is performed on these scanning artefacts corrected images using the MatchID software, employing parameters as given in Table 2, following procedures proposed by Vermeij et al. [32, 4]. Subsequently, the nanomechanical alignment framework [4] is applied to allow direct spatial correlation between EBSD and DIC data, as observed in the strain maps of all deformation steps in Figure 6(e), with the aligned grain boundary overlaid.
@@
Figure 6: (a1) Electron Backscatter Diffraction (EBSD) IPF map illustrating the microstructure of the Zn coating prior to deformation, with hexagonal shapes illustrating the crystal orientations. (a2) Zoom-in of IPF map around a tripple junction. (b) Scanning Electron Microscope (SEM) image, in in-lens Secondary Electron (SE) mode, displaying the Indium-Tin (InSn) speckle pattern used for Digital Image Correlation (DIC). The aligned grain boundaries are shown as overlay in red. (c) Effective strain field captured (inset of (e5) via DIC at a global strain of 0.059, showing crystal slip and grain boundary deformation. Areas surrounded by pink lines denote interpolated DIC data, predominantly on grain boundary locations. (d) Load-elongation curve from the in-situ tensile test with color-coded points corresponding to the strain levels in subfigures (e1)-(e5). (e1)-(e5) Effective strain fields at progressive tensile test increments, highlighting the evolution of strain localization and the alignment with EBSD data.
@@
At a first glance, the materials reveals a range of deformation mechanisms. Crystallographic slip can be observed within grains, both in the form of sharp and more diffuse slip bands. Additionally, localizations occur on top of grain boundaries, hinting at the occurrence of grain boundary sliding and/or migration. Figure 6(c) shows this more clearly with an inset of the strain field around a triple junction, for the final deformation step. Distinct slip, diffuse intragranular deformation and grain boundary activity can be observed. Note that the grain boundary deformation is very strong and resulted in a local degradation of the DIC pattern, such that interpolation of the DIC data was required (in the areas outlined with a thin pink line) [5]. However, the grain boundary deformation is outside the scope of this paper and will be subjected to identification by an alternative variation of the SSLIP method that is focused on the identification of grain boundary sliding [35]. Here, the goal is identification of the active slip system inside (all of) the grains.
@@
For each experimental case, we extract the deformation data of a single grain, allowed by the data alignment, and remove a band of 5 pixels along the grain boundary to avoid any grain boundary sliding and/or migration mechanisms. Additionally, the peak identification on the sinograms is performed to only retrieve peaks that are above 80 % of the maximum value are identified, with a maximum of 5 peaks.
@@
The first experimental grain that we investigate is shown in Figure 7. An impression of the crystallography and the deformation of the grain is given in Figure 7(a-b), consisting of the EBSD map and the effective strain field. At a first glance, slip occurs over only one specific slip plane, making it a relatively simple case to start the demonstration of the +SSLIP methodology.
@@
Figure 7: Results of slip system identification for a selected HCP Zn grain, showing one type of slip bands, using the +SSLIP method. (a) The EBSD map and rotated crystal shape illustrates the crystal orientation. (b) The effective strain field, $E_{eff}$, showing the deformation pattern. (c) The rotation field, $\gamma_{rot}$, resulting from rotation-correction by +SSLIP. (d) The residual strain fraction field, $Rf_{2D}^{L_{2}}$, resulting from +SSLIP, showing the fit between the experimental and theoretical displacement gradient tensor. Below that, the slip activity map of each slip system of each slip family (Basal, PrismI, PyrIA, PyrICA and PyrII) is shown. Maps with red crosses indicate that systems were not selected by Radon transform preselection. Each map also features the slip trace (red line) and direction (orange arrow), with the Schmid factor ($SF$) depicted as colored circles (colorbar below (a)). Note that the SF is not used in the +SSLIP method.
@@
Application of +SSLIP results in identification of the slip activity fields as shown in Figure 7. As expected, the Radon transform preselection only picks up one specific trace, which in this case can be attributed to 6 individual slip systems: all 3 Basal systems, 1 PyrIA system and 2 PyrICA systems. Among these, Basal-2 and Basal-3 are identified as the active slip systems, which is unsurprising since Basal slip is known to have the lowest CRSS in Zn [7, 36]. The rotation field in Figure 7(c), identified by +SSLIP , shows a negligible contribution for this grain. Additionally, the residual fraction in Figure 7(d) is rather low, indicating that the identification was successful. Note that the rotation field and the residual fraction field is set to NaN (i.e. shown as grey in the figures) when the residual fraction is above its threshold or when the effective strain is below its threshold.
@@
While this example may seem rather straightforward, the activities of the three Basal slip systems do not follow Schmid’s law globally. In fact, the lowest SF Basal system, Basal-3, shows most of the activity, while the highest SF system, Basal-1, shows no activity. Likely, the local loading conditions in the polycrystalline coating change the local stress state, rendering the SF insufficient to predict slip system activities. This example shows that it is insufficient to rely on the SF.
3.3 Conclusive vs. Inconclusive Identification
@@
As explained in Section 2.6, the identification can be inconclusive when slip systems have equal 2D (in-plane) kinematics, of which Figure 8 shows an example. The nondiscriminatory slip system groups, based on a 5angle threshold of the slip trace and project slip direction, are annotated by specific green symbols within the Schmid factor circles of the slip system activity fields. Note that, in general, the in-plane similarity of slip systems can be judged visually by assessing the difference between the plotted slip trace and direction. In this case, there are 3 groups of nondiscriminatory systems and slip systems PrimI-1 and PyrIA-2 are each part of two nondiscriminatory groups. The green rectangle shows that Basal-1 and PyrICA-2 are nondiscriminatory, but showing limited slip activities. The two other groups, annotated by the green ”+” sign (Basal-3, PrismI-1 and PyrIA-2) and circles (PrismI-1, PyrIA-1 and PyrIA-2) are very close to each other, but different enough to be separated into two groups. These two groups have the largest slip activities and it is likely that only Basal-3 is actually active, since it has a significant SF and low CRSS. Finally, this grain still has several unique slip systems, of which Basal-2 shows limited activity.
@@
Figure 8: Evaluation of slip system activities in an HCP Zn grain using the +SSLIP method, with a focus on nondiscriminatory slip systems. The EBSD map indicates the crystal orientation (a), with corresponding effective strain (b) and rotation fields (c) illustrating the deformation. The residual fraction field (d) demonstrates the method’s fit. Nondiscriminatory slip system groups are identified and marked with specific green symbols within the SF circles to denote their indistinguishability due to close angular proximity. Slip systems not preselected by the Radon transform are marked with red crosses.

### v2

@@ 2.3.1 SSLIP Analysis with Preselected Slip System Combinations
For each pair of slip systems, we now apply the SSLIP methodology, utilizing the optimization equation (Section 2.1) and also correcting for rotations, as explained below in Section 2.4. Subsequently, the selection of the optimal combination for each datapoint is governed by the residual fraction, $Rf_{2D}^{L_{2}}$, which must fall below a predefined threshold to ensure accurate identification. This residual fraction is calculated as
$Rf_{2D}^{L_{2}}=\frac{\left\|\mathbf{H}^{exp}-\mathbf{H}^{theo}\right\|_{2D}}{E_{eff}},$ (9)
where $\mathbf{H}^{exp}$ represents the experimental displacement gradient tensor, and $\mathbf{H}^{theo}$ denotes the theoretical displacement gradient tensor predicted by a combination of slip systems and rotation (see Section 2.4), and $E_{eff}$ is the effective strain, serving as a normalization factor.
The constraints imposed on potential slip system combinations are outlined in Table 2.
@@
Each combination is thereby assessed based on its residual fraction (Eq. 9), rotation angle and minimum equivalent strain, according to the values specified in Table 2. These values have been chosen as they were found to work robustly for the analysis of hundreds of Zn grains, of which a few challenging cases are discussed in Section 3. Among the potential combinations which satisfy these three criteria, the one with the lowest sum of slip magnitudes, thereby indicating the path of least resistance and conforming to the principle of minimum energy, is selected as the definitive combination of slip system activities for that particular datapoint.
@@
Since SSLIP will be applied with only two systems at a time, the optimization problem will have fewer degrees of freedom, which allows us to simultaneously identify the local elastic rotations. The inclusion of rotations was not needed with SSLIP for the less challenging cases of FCC and BCC, but is found to be necessary for HCP because of its large number of slip systems with linearly dependent in-plane deformation gradient tensors. In practice, these rotations can be caused by (full) grain rotation (e.g. due to grain boundary sliding), by compatibility effects near a grain boundary or by other microstructural features [4], or simply by misalignment of the sample between (quasi in-situ) measurement steps. Therefore, we add (local) rotation correction to the methodology, which will not only improve the slip system identification but will also give an extra result, i.e. the rotation field, which may be useful for further analysis.
The correction involves incorporating an additional rotation component into SSLIP, represented by an approximated rotation tensor. For small rotations, which are common in our observations, the in-plane rotation tensor can be approximated as follows:
@@
Here, $\theta$ represents the rotation angle and, under the assumption that $\theta$ is small, $\sin(\theta)\approx\theta$ and $\cos(\theta)\approx 1$, allowing us to capture the in-plane rotation with this approximated matrix, in displacement gradient tensor form. It is then incorporated into SSLIP as an additional ”slip system”, for each combination that will be assessed. This results in the rotation value $\theta$ at every datapoint, along with the slip system magnitudes. In cases where the rotations would be so large that the simplified rotation tensor is a poor estimate, the full version of $\mathbf{H}^{rot}$ in Eq. 10 can be used, at the expense that $\mathbf{H}^{rot}$ needs to be updated using the current estimate of the rotation angle, at each location and at each step in the iterative optimization routine. After successful identification, the 2D kinematics at each data point is described as
$\mathbf{H}^{theor}=\gamma^{i}\mathbf{P}^{i}_{2D}+\gamma^{j}\mathbf{P}^{j}_{2D}+\mathbf{H}^{rot},$ (11)
in which $i$ and $j$ are the two identified slip systems.
@@ 2.5 Validation on Virtual Experiment
Now, the full +SSLIP methodology is applied to the challenging HCP virtual experiment, for which the results are shown in Figure 5. This figure is structured identically to those in the experimental case studies below. Figure 5(a) shows the crystallographic orientation of the virtual experiment and Figure 5(b) shows the effective strain field. Application of the Radon transform results in eight preselected slip systems, see Figure 2. Subsequently, application of Rotation-corrected SSLIP on all pairs of the slip systems ($N_{duo}=28$) and point-wise selection of the pair with the lowest residual, results in the slip activity fields as shown in Figure 5. The correct slip systems were Basal-3, PrismI-3, PyrICA-9 and PyrII-5, all of which show a single slip band in their respective slip activity field, oriented as expected along their slip trace. The other four preselected but non-active slip systems show no activity over the full area, as expected. The error of the SSLIP analysis is represented by the residual norm fraction in Figure 5(d), $Rf_{2D,max}^{L_{2}}$ (Eq. 9), for which low values are observed. At the edges of the slip bands, the errors are higher because the error is normalized with the effective strain, which is low in these areas, leading to a lower signal to noise ratio. No rotation was applied in the virtual experiment, which is reproduced by the +SSLIP algorithm as shown in Figure 5(c). The intersection of three slip bands in the center of the domain cannot be identified (grey area), since only two slip systems can be active per datapoint. Importantly, these intersections are also not falsely attributed to incorrect slip systems. In practice, in the unlikely case when more than two systems overlap, the slip activity at such multi-slip intersections can easily be inferred from the neighbouring slip activity. Overall, the identification shown here is robust and accurate and shows promise for identification of complex HCP activities.
@@
Figure 5: Demonstration of the +SSLIP results on the challenging virtual experiment. (a) The EBSD map and crystal shape illustrate the crystal orientation of this virtual experiment. (b) The effective strain field, $E_{eff}$, clearly showing four slip bands. (c) The rotation field, $\gamma_{rot}$, resulting from rotation-correction by +SSLIP . (d) The residual strain fraction field, $Rf_{2D}^{L_{2}}$, resulting from +SSLIP, showing the fit between the ”experimental” and theoretical displacement gradient tensor. Below that, the slip system activity map of each slip system of each slip family (Basal, PrismI, PyrICA and PyrII) is shown. Maps with red crosses indicate that systems were not selected by Radon transform preselection. Each map also features the slip trace (red line) and direction (orange arrow), with the Schmid factor ($SF$) depicted as colored circles (colorbar below (a)). The SF is not used in the +SSLIP method, but plotted here such that structure of the figure corresponds to the experimental case studies
2.6 Conclusive vs. Inconclusive Identification
@@
The +SSLIP method is demonstrated to accurately identify slip system activities on the virtual experiment, however, when many slip systems are considered, the kinematics of certain combinations of systems may not be distinguishable in 2D (i.e. in-plane). In theory, even though each slip system has a unique 3D displacement gradient, the 2D part of the displacement gradient tensor could be exactly, or almost, the same for two (or more) slip systems, making those slip systems fundamentally nondiscriminatory in the SSLIP method. In practice, however, the minimization of the sum of slip activities, as used in the original SSLIP method [22], was found to be robust against slip systems with almost equal 2D displacement gradient tensors, therefore, this issue was not considered in detail before. For HCP, however, there are many potential slip systems from multiple different families, which typically cannot be excluded a-priori due to unknown Critical Resolved Shear Stress (CRSS) values. As a result, the occurrence of multiple highly similar 2D displacement gradient tensors was found to be problematic for certain grain orientations, a few examples of which are treated in Section 3.3. In such cases, a group of two or more slip systems can be defined with similar enough 2D displacement gradient tensors such that the measured slip activity can only be assigned to the group of slip systems. When slip activity is assigned to such a ’nondiscriminatory slip system group’, we call the identification ’inconclusive’, as counterpart to the regular ’conclusive’ identification where the slip activity is uniquely assigned to one or more slip systems.
@@
Therefore, in the +SSLIP method, after the Radon transform preselection, all possible nondiscriminatory slip system groups are identified from the available preselected slip systems, before executing the +SSLIP identification. The likelihood of misidentification due to the existence of slip systems that are similar in their 2D kinematics, depends on the absolute measurement accuracy of the slip trace angle and the projected slip direction angle. In turn, these depend on the combined accuracy of EBSD, SEM-DIC (including SEM scanning artefacts), and their mutual alignment. In this work, it was found that these two angles can have an error up to $\sim$5 . Therefore, we define an angle threshold of 5 , i.e., if two preselected slip systems have both a slip trace angle and projected slip direction angle within 5 of each other, they are considered nondiscriminatory. Through this procedure, the preselected slip systems are, when necessary, included in nondiscriminatory slip system groups in the +SSLIP identification. These groups can consist of more than two slip systems. In the (pixel-wise) +SSLIP output, the slip activity assigned to a nondiscriminatory group is attributed equally to each slip system within the group. The way this works in practice will be shown in the section below (note that the virtual test case in Figure 5 is a regular example of a ’conclusive’ identification).
@@ 3 Experimental Case Study: +SSLIP applied on HCP Zinc Coating
In this section, we demonstrate the strength of the +SSLIP method on experimental SEM-DIC results of an HCP Zinc anti-corrosion coating, which will show activities from multiple slip families: Basal, PyrI and PyrII. The focus will be on challenging case studies that include multiple ’nondiscriminatory slip system groups’. Additionally, we will showcase robust identification on grains that exhibit cross slip and diffuse slip, for which pure trace-based analysis is hardly possible [22].
In Section 3.1, we briefly detail the experimental methodology for acquiring microstructure-aligned SEM-DIC strain fields. Section 3.2 demonstrates +SSLIP on a simple case of single slip and Section 3.3 shows how an inconclusive identification can be interpreted. More complex cases of (diffuse) multi-slip and cross slip are treated in Sections 3.4 and 3.5.
@@ 3.1 Material Characterization and SEM-DIC Testing
The material used in this case study is a hot-dip galvanized skin-passed Zn coated steel sample with material code DX54-Galvanized Iron (GI). Samples are cut using wire-EDM (gauge width 4 mm, thickness 0.7 mm). Before testing, the sample surface is carefully prepared in a single step of polishing using Oxide Polishing Suspension (OPS)-NonDry and ethanol (1:1) suspension for 10 minutes. This short polishing provides the optimal compromise between good surface quality and low material removal, since the Zn coating is only $\sim 10\mu m$.
@@
Next, several $\sim$50 $\mu$m sized Regions of Interest (ROIs) are chosen for characterization and in-situ SEM-DIC tensile testing, as shown in Figure 6 for one ROI. EBSD is performed with an EDAX Digiview 2 camera, with offline spherical indexing using EMSphInx for improved quality [33], resulting in an Inverse Pole Figure (IPF) map as visualized in Figure 6(a). Subsequently, we apply a fine and dense InSn SEM-DIC speckle pattern applied according to the parameters of pattern b in Table 1 of Hoefnagels et al. [34], resulting in approximately 90-100 nm sized speckles, see Figure 6(b). The specimen is then mounted in a Kammrath&Weiss micro-tensile stage, which is installed inside a Tescan Mira 3 SEM for interrupted in-situ testing, at deformation steps as shown in Figure 6(d). For SEM-DIC imaging, we use in-lens SE imaging (5 kV beam voltage at 7 mm working distance) [35, 4, 5]. At each deformation step, a horizontal and vertical scan are captured that are combined into a single image, thereby correcting for SEM scanning artefacts, using ScanCorr [4]. Subsequently, DIC is performed on these corrected images using MatchID DIC software, employing parameters as given in Table 3, following procedures proposed by Vermeij et al. [35, 4]. Subsequently, a nanomechanical alignment framework [4] is applied to allow direct spatial correlation between EBSD and DIC data, as observed in the strain maps of all deformation steps in Figure 6(e), with the aligned grain boundaries overlaid.
@@
Figure 6: (a1) EBSD (out-of-plane) IPF map illustrating the microstructure of the Zn coating prior to deformation, with hexagonal shapes illustrating the crystal orientations. (a2) Zoom-in of IPF map around a triple junction. (b) Scanning Electron Microscope (SEM) in-lens Secondary Electron (SE) scan displaying the Indium-Tin (InSn) speckle pattern used for Digital Image Correlation (DIC). The aligned grain boundaries are shown as overlay in red. (c) Effective strain field captured (inset of (e5) via DIC at a global strain of 0.059, showing crystallographic slip and grain boundary deformation. Areas surrounded by pink lines denote interpolated DIC data, predominantly at grain boundary locations. (d) Load-elongation curve from the in-situ tensile test with color-coded points corresponding to the strain levels in subfigures (e1)-(e5). (e1)-(e5) Effective strain fields at progressive tensile test increments, highlighting the evolution of strain localization and the alignment with EBSD data.
@@
At first glance, the material reveals a range of deformation mechanisms. Crystallographic slip can be observed within grains, both in the form of sharp and more diffuse slip bands. Additionally, localizations occur on top of grain boundaries, hinting at the occurrence of grain boundary sliding and/or migration. Figure 6(c) shows this more clearly with an inset of the strain field around a triple junction, for the final deformation step. Discrete slip, diffuse intragranular deformation, and grain boundary activity can be observed. Note that the grain boundary deformation is very strong and resulted in a local degradation of the DIC pattern, such that interpolation of the DIC data was required (in the areas outlined with a thin pink line), as described in [5]. Grain boundary deformation is outside the scope of this paper and will be subjected to identification by an alternative variation of the SSLIP method focused on grain boundary sliding [36].
@@
The first experimental grain that we investigate is shown in Figure 7. The crystallography and the deformation of the grain (at the final deformation increment) is given in Figure 7(a-b). At a first glance, slip occurs over only one specific slip plane, making it a relatively simple case to start the demonstration of the +SSLIP methodology.
@@
Figure 7: Results of slip system identification for a selected HCP Zn grain, showing single slip, using the +SSLIP method. (a) The EBSD map and rotated crystal shape illustrates the crystal orientation. (b) The effective strain field, $E_{eff}$, showing the deformation pattern. (c) The rotation field, $\gamma_{rot}$, resulting from rotation-correction by +SSLIP . (d) The residual strain fraction field, $Rf_{2D}^{L_{2}}$, resulting from +SSLIP , showing the fit between the experimental and theoretical displacement gradient tensor. Below that, the slip activity map of each slip system of each slip family (Basal, PrismI, PyrIA, PyrICA and PyrII) is shown. Maps with red crosses indicate that systems were not selected by Radon transform preselection. Each map also features the slip trace (red line) and direction (orange arrow), with the SF depicted as colored circles (colorbar below (a)). Note that the SF is not used in the +SSLIP method.
@@
Application of +SSLIP results in identification of the slip activity fields as shown in Figure 7. As expected, the Radon transform preselection only picks up one specific slip band, which in this case matches with six individual slip systems: all 3 Basal systems, 1 PyrIA system and 2 PyrICA systems. Among these, Basal-2 and Basal-3 are identified as the active slip systems, which is unsurprising since Basal slip is known to have the lowest CRSS in Zn [7, 37]. The rotation field in Figure 7(c), as identified by +SSLIP , shows a negligible contribution for this grain. Additionally, the residual fraction in Figure 7(d) is low and shows no significant signs of slip bands, indicating that the identification was successful. Note that the rotation field and the residual fraction field is set to NaN (shown as grey in the figures) when the residual fraction is above its threshold or when the effective strain is below its threshold.
@@
While this example may seem rather straightforward, the activities of the three Basal slip systems do not follow Schmid’s law globally. In fact, the lowest SF Basal system, Basal-3 (SF of 0.07), shows most of the activity, while the highest SF system, Basal-1 (SF of 0.23), shows no activity. This could be due to the fact that the local loading conditions in the polycrystalline coating change the local stress state, rendering the SF insufficient to predict slip system activities. However, definite conclusions require measurements of the local stress state or advanced Transmission Electron Microscopy (TEM) analysis of dislocations, which is outside the scope of this work.
3.3 Conclusive vs. Inconclusive Identification results
@@
As explained in Section 2.6, the identification can be inconclusive when slip systems have equal 2D (in-plane) kinematics, of which Figure 8 shows an example. The nondiscriminatory slip system groups, based on a 5 angle threshold of the slip trace and projected slip direction, are annotated by specific green symbols within the SF circles of the slip system activity fields. Note that, in general, the in-plane similarity of slip systems can be judged visually by assessing the difference between the plotted slip trace and direction. In this case, there are 3 groups of nondiscriminatory systems. Slip systems PrismI-1 and PyrIA-2 are each part of two nondiscriminatory groups. The green rectangle shows that Basal-1 and PyrICA-2 are nondiscriminatory, but showing limited slip activities. The two other groups, annotated by the green ”+” sign (Basal-3, PrismI-1 and PyrIA-2) and circles (PrismI-1, PyrIA-1 and PyrIA-2) are very close to each other, but different enough to be separated into two groups. These two groups have the largest slip activities and it is likely that only Basal-3 is actually active, since it has a significant SF ($SF>0.35$) and it is expected to have a low CRSS. Finally, this grain still has several unique slip systems, of which Basal-2 shows limited activity.
@@
Figure 8: Evaluation of slip system activities in an HCP Zn grain using the +SSLIP method, with a focus on nondiscriminatory slip systems. See the caption of Figure 7 for the figure structure. Nondiscriminatory slip system groups are identified and marked with specific green symbols within the SF circles to denote their indistinguishability due to close angular proximity.